The Lakers scored 43 points on a combination of 20 two point baskets and three point baskets. How many shots of each type were made?
use x to represent two point baskets and y to rpresent three point baskets. Write an equation that ilustrates the total number of baskets made during the game.
Write an equation that illustrates how many points total were made during the game.
Use the equations from to solve this exercise as a system of equations. Use substitution or elimination to determine how many shots of each type were made.

Answer by dfrazzetto(283)


x+y=20
2x+3y=43 multiply top equation by -2
-2x+-2y = -40
2x+ 3y = 43
------------------
y = 3
x+y=20
x+3=20
x=20-3 = 17
So, 17 two point baskets and 3 three point baskets
CHECK:
x=17, y=3
2x+3y=43
2(17)+3(3) = 43
34 + 9 = 43
43=43 √